OpenStudy (raffle_snaffle):

Given a damped mass system in which the mass is subject to harmonic forcing of Fo*sin(wt), show that the peak amplitude of the vibration occurs at a frequency wp given by the expression. wp/wn = [1 - 2(Z^2)]^(1/2) where Z is the damping ratio

OpenStudy (raffle_snaffle):

@IrishBoy123

OpenStudy (raffle_snaffle):

wp = peak frequency

OpenStudy (raffle_snaffle):

wn = natural frequency

OpenStudy (irishboy123):

yeah, i got it!! so we need to solve a 2nd order non-hom DE it's all been transcribed on Wiki , thankfully

OpenStudy (raffle_snaffle):

I am going to upload a picture of my notes. I want to know if we apply these steps.

OpenStudy (irishboy123):

sure but you know what the story is here? you're solving a 2nd order non-hom DE ?

OpenStudy (irishboy123):

and then doing a load of really shιtty algebra to match stuff up

OpenStudy (raffle_snaffle):

So the homogenous part goes to zero and we just take the derivative of the particular solution.

OpenStudy (irishboy123):

you will always have a complementary solution and a particular solution.

OpenStudy (raffle_snaffle):

For a long period of time, the complementary goes to zero though.

OpenStudy (irishboy123):

no! the complimentary is the perpetual oscillation for a spring-mass system. the eternal exchange of energy. it is then conditioned by the damping term. which is the loss of energy

OpenStudy (irishboy123):

which can also be included in the complementary....so you are right

OpenStudy (raffle_snaffle):

Okay.

OpenStudy (raffle_snaffle):

The natural frequency goes to zero.

OpenStudy (raffle_snaffle):

I am working on it now. I will show you my solution when I am done.

OpenStudy (raffle_snaffle):

I will post it up. I will brb and then tag you.

OpenStudy (irishboy123):

i will try work a solution too

OpenStudy (raffle_snaffle):

What I have so far.

OpenStudy (irishboy123):

Your stuff looks good and ends I think here: \(x(t) = - \dfrac{F_o c \omega}{ ( c \omega )^2 + (k - m \omega^2)^2} \cos \omega t + \dfrac{F_o (k - m \omega^2))}{(c \omega )^2 + (k - m \omega^2)^2} \sin \omega t \) if we replace with \(z = \dfrac{c}{2m}\) and \(\omega_n = \sqrt{\dfrac{k}{m}}\) it becomes \(x(t) = - \dfrac{F_o 2mz \omega}{ ( 2mz \omega )^2 + (m \omega_n^2 - m \omega^2)^2} \cos \omega t + \dfrac{F_o (m \omega_n^2 - m \omega^2))}{(2mz \omega )^2 + (m \omega_n^2 - m \omega^2)^2} \sin \omega t \) and \(x(t) = - \dfrac{2 z \omega F_o /m}{ ( 2z \omega )^2 + ( \omega_n^2 - \omega^2)^2} \cos \omega t + \dfrac{ ( \omega_n^2 - \omega^2) F_o/m }{(2z \omega )^2 + ( \omega_n^2 - \omega^2)^2} \sin \omega t \) In this form it is easy to turn into a single sinusoid using an addition formula \(\sin (A - B) = \sin A \cos B - \sin B \cos A\) Looks more obvious this way.... \(x(t) =\dfrac{F_o}{m} \dfrac{1}{ \sqrt {( 2z \omega )^2 + ( \omega_n^2 - \omega^2)^2}} \left( - \dfrac{2 z \omega }{\sqrt{ ( 2z \omega )^2 + ( \omega_n^2 - \omega^2)^2}} \cos \omega t \\ + \dfrac{ ( \omega_n^2 - \omega^2) }{\sqrt { (2z \omega )^2 + ( \omega_n^2 - \omega^2)^2}} \sin \omega t \right)\) Making it: \(x_p=\dfrac{F_o}{m} \dfrac{1}{ \sqrt {( 2z \omega )^2 + ( \omega_n^2 - \omega^2)^2}} \sin (\omega t - \psi) \qquad \star\) That way you can read the amplitude straight off. If you differentiate the non-sinusoid bit of \(\star\) wrt \(\omega\) you should then get \(\omega\) for the max amplitude of the steady state oscillation. Without some IV's I don't see how you can say much about the transient decaying oscillation that is your complimentary solution. Which again I would also be tempted to write in this combined form as \(x_c = A e^{- zt } \sin (\omega_n t + \phi)\) Furthermore I get a slightly different answer to the one you are looking for so maybe I am goofing along the way.

OpenStudy (raffle_snaffle):

Guess what... All she wanted was to show a proof... Lol

OpenStudy (irishboy123):

yes but there's no proof or her equation is wrong or i'm a fucκwit. most likely outcome.